Solve: 8y^3 + 27 divided by 2xy - 10y + 3x -15
Answer by richwmiller(17219)

Solve: 8y^3 + 27 divided by 2xy - 10y + 3x -15
This is not an equation and so can't be solved. It is an expression and can be reduced or simplified.
8y^3 + 27=(2y+3)*(4y^2-6y+9)
2xy - 10y + 3x -15=(x-5)*(2 y+3)
(2y+3)*(4y^2-6y+9)/(x-5)*(2 y+3)
(4y^2-6y+9)/(x-5)
This is the furthest this can go since we have x in the denominator and and y in the numerator